### v1

# Generalist World Model Pre-Training for Efficient Reinforcement Learning

Abstract
Sample-efficient robot learning is a longstanding goal in robotics. Inspired by the success of scaling in vision and language, the robotics community is now investigating large-scale offline datasets for robot learning. However, existing methods often require expert and/or reward-labeled task-specific data, which can be costly and limit their application in practice. In this paper, we consider a more realistic setting where the offline data consists of reward-free and non-expert multi-embodiment offline data. We show that generalist world model pre-training (WPT), together with retrieval-based experience rehearsal and execution guidance, enables efficient reinforcement learning (RL) and fast task adaptation with such non-curated data. In experiments over 72 visuomotor tasks, spanning 6 different embodiments, covering hard exploration, complex dynamics, and various visual properties, WPT achieves 35.65% and 35% higher aggregated score compared to widely used learning-from-scratch baselines, respectively.
Reinforcement Learning, World Models, Model-based Reinforcement Learning, Continuous Control, Machine Learning
1 Introduction
Figure 1: Aggregate performance on 50 manipulation tasks from Meta-World and 22 locomotion tasks from DMControl with pixel inputs. We pre-train one single world model per benchmark on non-curated offline data, which can be fine-tuned to largely improve RL’s sample efficiency. With 150k online samples, WPT obtains 35.65% and 35% higher normalized scores compared to DrQ v2 and Dreamer v3 using the same samples and matches baselines’ results obtained with higher sample budgets (500k samples for DMControl and 1M samples for Meta-World).
Figure 2: Overview of generalist world model pre-training (WPT). Facing non-curated offline datasets with reward-free, non-expert, and multi-embodiment data, we train a task and embodiment-agnostic world model. The pre-trained world model improves the sample efficiency of RL training over a wide range of tasks.
With the success of scaling laws in vision and language (Brown et al., 2020; He et al., 2022; Kirillov et al., 2023; Touvron et al., 2023), the robotics community has started investigating the use of large-scale offline datasets for robot learning (Brohan et al., 2023b; Walke et al., 2023; Brohan et al., 2023a; O’Neill et al., 2023; Khazatsky et al., 2024; Team et al., 2024). Large-scale offline datasets enable the trained agents to learn a wide range of skills as well as help the agent adapt to new tasks with limited samples (Team et al., 2024).
To leverage offline datasets for decision-making, imitation learning (IL) (Brohan et al., 2023b; Team et al., 2024; Doshi et al., 2024) and offline RL (Levine et al., 2020; Kumar et al., 2020; Fujimoto & Gu, 2021; Kumar et al., 2023) are commonly used. However, these methods have limitations in terms of both data and algorithmic aspects. IL assumes the availability of expert data from human demonstration or specialist RL agents, which usually involves an expensive data collection procedure. Further to this, agents learned by IL are usually sensitive to environmental perturbations such as action delay, lighting conditions, sensor noise, etc (Chae et al., 2022; Zare et al., 2024). Offline RL aims to learn agents from offline data but suffers from training instability, especially for pixel-based observations (Kumar et al., 2020; Lu et al., 2023). Besides the training stability issue, offline RL requires task-specific datasets, labeled with rewards. This poses another challenge for many real-world applications, e.g., robotic manipulation, to utilize offline datasets for new tasks, it is necessary to retrospectively annotate the image-based offline data with rewards, which can be challenging and laborious. Furthermore, fine-tuning an offline RL agent, so-called offline-to-online RL (Nair et al., 2020; Lee et al., 2022; Zhao et al., 2022; Yu & Zhang, 2023; Nakamoto et al., 2024), requires additional tricks to stabilize training and prevent performance collapse caused by the distributional shift in the fine-tuning stage.
In contrast to IL and offline RL, which require curated datasets, many real-world scenarios have data from various sources. To this end, we focus on the more realistic setting where the offline dataset includes non-curated mixed-quality data consisting of reward-free and non-expert data, collected by multiple agents with varying embodiments. This setting has a minimal requirement for offline data, which significantly expands the pool of usable data. The primary research question we aim to answer is:
What is the best way to leverage non-curated offline data for efficient multi-task & multi-embodiment learning?
Definition 1.1 (Non-curated data).
Non-curated data encapsulates all observation-action trajectories, i.e., it can consist of reward-free and non-expert data for any embodiment.
Existing methods for leveraging offline data (e.g. IL and offline RL) typically fail given non-curated data. For example, IL would require manually filtering the dataset to select expert trajectories and offline RL would require retrospective reward labeling. With such non-curated data, a common strategy is to pre-train a visual encoder (Schwarzer et al., 2021; Nair et al., 2022; Parisi et al., 2022; Xiao et al., 2022; Yang & Nachum, 2021; Shang et al., 2024). However, visual pre-training alone fails to fully leverage the rich information that exists in the offline dataset, e.g., dynamics model, informative states, or action prior, etc.
Facing the non-curated offline dataset, training a world model seems a “natural” choice, as it holds the promise of better generalization and better sample efficiency. However, model-based approaches are not commonly used in the offline-to-online setting (Rafailov et al., 2023; Wu et al., 2024). This is partially due to the performance margin between model-based and model-free approaches not being significant (Yu et al., 2020; Kidambi et al., 2020) while model-based approaches involve higher training complexity. In addition, in the typical offline-to-online RL setting, training an offline agent with model-based approaches can be unstable, especially for pixel inputs (Lu et al., 2023), and strong regularization is required during the online training stage (Rafailov et al., 2023).
In this paper, we show that with careful design choices, world model pre-training unlocks more useable offline data and better fits the pre-train and fine-tune paradigm used for scaling. Our key insights are i) instead of pre-training a policy using model-based offline RL as done in previous work (Yu et al., 2020; Kidambi et al., 2020; Rafailov et al., 2023), learning a world model is stable and scalable, ii) unlike previous methods (Hafner et al., 2020, 2021, 2023; Hansen et al., 2024) which learns a world model per task, by simply padding actions to unify the action spaces, a single world model can be pre-trained on the non-curated offline data consisting of data sources from different tasks and different embodiments, iii) via fine-tuning, a pre-trained generalist world model can boost RL’s sample efficiency over a wide range of tasks and embodiments, iv) during fine-tuning, experience rehearsal, and execution guidance can largely improve the performance.
Table 1: Comparison with different policy learning methods that leverage offline data.
Offline RL Off2On RL RLPD MT IL MT Offline RL WPT (ours)
Reward-free offline data
Non-expert offline data
X-embodiment offline data
Continual Improvement
Training stability
We extensively evaluate WPT on 72 pixel-based continuous control tasks covering locomotion and manipulation with different action spaces, hard exploration, high dimensions, and complex dynamics. Under a limited sample budget (150k samples), WPT outperforms previous SOTA methods by a decent margin. Specifically, WPT obtains 35.65% and 35% higher normalized scores compared to DrQ v2 and Dreamer v3 under the same sample budget and matches baselines’ results obtained with higher sample budgets (500k samples for DMControl and 1M samples for Meta-World.). For example, WPT enables an agent learning to control an Ant robot to walk forward within 100 trails, while widely used learning-from-scratch baselines need 10-30$\times$ samples. We further demonstrate that, without any modifications, WPT also helps task adaptation, where the agent is required to continually adapt its skill to new tasks. Our results demonstrate strong performance with a large margin over model-free methods when leveraging non-curated offline data. We hope that our results motivate further investigation of model-based approaches for using offline data. In summary, our contributions include:
  • We propose a new and more realistic setting for leveraging offline data where the offline data consists only of reward-free and non-expert multi-embodiment data.
  • We show that generalist world model pre-training along with a series of careful design choices achieves strong performance in 72 visuomotor tasks spanning 6 embodiments.
  • To facilitate further research with non-curated offline data, we open-source our datasets, including 60K trajectories and 10M transitions from the DMControl and Meta-World benchmarks.
2 Related Work
In this section, we review methods which leverage offline data, including pre-training in the context of RL and the so-called generalist agents. See Table 1 for a comparison of what types of data the methods can use.
RL with task-specific offline datasets   Leveraging offline data is a promising direction to improve the sample efficiency in RL. One representative method is Offline RL. Offline RL trains agents from offline data without interacting with the environment. It usually constrains the distance between the learned policy and behavior policies in different ways (Kumar et al., 2019, 2020; Wu et al., 2019; Kostrikov et al., 2021, 2022; Fujimoto & Gu, 2021; Uchendu et al., 2023). However, the performance of the learned policy is highly affected by the quality of offline datasets (Yarats et al., 2022). To continue improving the policy, offline-to-online RL (Nair et al., 2020; Zhao et al., 2022; Lee et al., 2022; Yu & Zhang, 2023; Rafailov et al., 2023) was proposed.
Offline RL and offline-to-online RL commonly suffer from training instability and require additional treatments (Lee et al., 2022; Lu et al., 2023). RLPD (Ball et al., 2023) shows that off-policy RL delivers strong performance by directly concatenating the offline data with online data. However, RLPD still requires reward-labeled task-specific offline data and hasn’t discussed the multi-embodiment cases. Recently, ExPLORe (Li et al., 2023) labels reward-free offline data with approximated upper confidence bound (UCB) of rewards to solve hard exploration tasks. However, ExPLORe obtains good performance when the offline data includes near-expert data of the target tasks, while we consider a more general setting with non-curated offline data.
RL with multi-task offline datasets   Recently, several works have applied offline RL in a multi-task setting (Kalashnikov et al., 2021; Yu et al., 2021; Kumar et al., 2023; Julian et al., 2020; Hansen et al., 2024), but require known rewards.  Georgiev et al. (2024) pre-trains a world model for multi-task RL but is designed for state-based inputs and requires reward-labeled offline data for world model pre-training.
To lift the assumption of known rewards, human labeling (Cabi et al., 2020; Singh et al., 2019), inverse RL (Ng et al., 2000; Abbeel & Ng, 2004) or generative adversarial imitation learning (Ho & Ermon, 2016) can be applied. However, this requires human labor or expert demonstrations. Yu et al. (2022) directly sets zero rewards for unlabelled data, which introduces additional bias.
Another stream of work leverages in-the-wild data for RL training. Most methods focus on representation learning. Stooke et al. (2021); Yang & Nachum (2021); Schwarzer et al. (2021); Shah & Kumar (2021); Yuan et al. (2022); Wang et al. (2022); Parisi et al. (2022); Sun et al. (2023); Ze et al. (2023) investigate different representation learning methods and show promising performance boosts during fine-tuning.
Generalist Agents   RL methods usually perform well on a single task (Vinyals et al., 2019; Andrychowicz et al., 2020), however, this contrasts with humans that can perform multiple tasks well. Recent works have proposed generalist agents that master a diverse set of tasks with a single agent (Reed et al., 2022; Brohan et al., 2023b; Team et al., 2024). These methods typically resort to scalable models and large datasets and are trained via imitation learning (Brohan et al., 2023b, a; O’Neill et al., 2023; Khazatsky et al., 2024). In contrast, we seek to train a generalist world model and use it to boost RL performance for multiple tasks and embodiments.
World models   World models learn to predict future observations or states based on historical information. World models have been widely investigated in online model-based RL (Ha & Schmidhuber, 2018; Hafner et al., 2020; Micheli et al., 2023; Alonso et al., 2024). Recently, the community has started investigating scaling world models (Ha & Schmidhuber, 2018), for example, Hu et al. (2023a); Pearce et al. (2024); Wu et al. (2025); Agarwal et al. (2025) train world model with Diffusion Models or Transformers. However, these models are usually trained on demonstration data. In contrast, we explore the offline-to-online RL setting – closely fitting the successful pre-train and then fine-tune paradigm – and we focus on leveraging reward-free and non-expert data to increase the amount of available data for pre-training.
3 Methods
In this section, we detail our two-stage approach, which consists of (i) world model pre-training, which learns a generalist (i.e. multi-task & embodiment) world model, given offline data, which rather importantly, includes reward-free and non-expert data, and (ii) RL-based fine-tuning which leverages the pre-trained world model and online interaction in an offline-to-online fashion. See Fig. 2 and Alg. 1.
3.1 Problem Setup
In this paper, we assume the agent can access a static offline dataset $\mathcal{D}_{\text{off}}$, which consists of trajectories $\{\tau_{\text{off}}^{i}\}_{i=1}^{N_{\text{off}}}$, where $\tau_{\text{off}}^{i}$ includes observations and actions $\{o^{i}_{t},a^{i}_{t}\}_{t=0}^{T}$ collected by unknown behavior policies. This means that (i) rewards $r^{i}_{t}$ are unknown, (ii) $\mathcal{D}_{\text{off}}$ does not necessarily include expert trajectories, and (iii) datasets consist of multi-embodiment data. The agent interacts with the environment to collect labeled trajectories $\tau_{\text{on}}^{i}=\{o^{i}_{t},a^{i}_{t},r^{i}_{t}\}_{t=0}^{T}$ and stores them in an online dataset $\mathcal{D}_{\text{on}}=\{\tau_{\text{on}}^{i}\}_{i=1}^{N_{\text{on}}}$. The goal of this paper is to learn a high-performance policy from both $\mathcal{D}_{\text{off}}$ and $\mathcal{D}_{\text{on}}$ whilst minimizing the amount of online interaction ($N_{on}$) by leveraging the unlabeled offline data $\mathcal{D}_{\text{off}}$. Removing the requirement of rewards and expert trajectories enables the agent to leverage a large set of diverse datasets.
3.2 Multi-Embodiment World Model Pre-training
During pre-training, instead of training one model per task – as done in previous work (Hafner et al., 2020, 2021, 2023; Hansen et al., 2024) – we show that a single multi-task & embodiment world model more effectively leverages the offline data. Compared to methods (Yang & Nachum, 2021; Schwarzer et al., 2021; Yuan et al., 2022; Parisi et al., 2022; Ze et al., 2023) that only pre-train representations, pre-training a world model creates a general understanding of the environment, which we show can further boost agents’ performance during fine-tuning.
We adopt a widely used world model in the literature, the recurrent state space model (RSSM) (Hafner et al., 2019) by making several modifications: (i) we remove the task-related losses, (ii) we pad the action with zeros to unify the action dimension of different embodiments, and (iii) we scale the model size to 280M. Given these modifications, we show that RSSMs can successfully learn the dynamics of multiple embodiments and can be fine-tuned for mastering different tasks. Our first stage consists of pre-training the following components:
Sequence model $:h_{t}=f_{\theta}(h_{t-1},z_{t-1},a_{t-1})$
Encoder $:z_{t}\sim q_{\theta}(z_{t}|h_{t},o_{t})$
Dynamics predictor $:\hat{z}_{t}\sim p_{\theta}(z_{t}|h_{t})$
Decoder $:\hat{o}_{t}\sim d_{\theta}(\hat{o}_{t}|h_{t},z_{t}).$
The models $f_{\theta}$, $q_{\theta}$, $p_{\theta}$ and $d_{\theta}$ are optimized jointly by minimizing:
$\mathcal{L}(\theta)=\mathbb{E}_{p_{\theta},q_{\theta}}\Big{[}\sum_{t=1}^{T}\underbrace{-\ln p_{\theta}(o_{t}|z_{t},h_{t})}_{\text{pixel reconstruction loss}}\\ +\beta\cdot\underbrace{\text{KL}\big{(}q_{\theta}(z_{t}|h_{t},o_{t})\ ||\ p_{\theta}(z_{t}|h_{t})\big{)}}_{\text{latent state consistency loss}}\Big{]}.$ (1)
The first term minimizes the reconstruction error while the second term enables the latent dynamics learning. Note that we removed the task-related objectives, e.g., reward prediction and continue prediction. We note that there is plenty of room to improve the world model’s pre-training, e.g., by leveraging recently developed self-supervised training methods (Eysenbach et al., 2023), or advanced architectures (Vaswani et al., 2017; Gu et al., 2022). These improvements are orthogonal to our method and we leave it as future work.
3.3 RL-based Fine-tuning
In our fine-tuning stage, we allow the agent to interact with the environment to collect new data $\tau_{\text{on}}^{i}=\{o^{i}_{t},a^{i}_{t},r^{i}_{t}\}_{t=0}^{T}$. The newly collected data is used to learn a reward function via supervised learning whilst also fine-tuning the world model with Eq. 1. For simplicity, we represent the concatenate of $h_{t}$ and $z_{t}$ as $s_{t}=[h_{t},z_{t}]$. The actor and critic are trained with imagined trajectories $\tilde{\tau}(s,a)$ generated by rolling out the policy $\pi_{\phi}(a|s)$ in the world model $p_{\theta}$ starting from the initial state distribution $p_{0}(s)$, which are samples from replay buffer. Following DreamerV3, the critic $v_{\phi}(V_{t}^{\lambda}\mid s_{t})$ learns to approximate the distribution
over the $\lambda$-return $V_{t}^{\lambda}$ calculated as,
$\underbrace{V^{\lambda}_{t}}_{\lambda-\text{return}}=\hat{r}_{t}+\gamma~{}\begin{cases}(1-\lambda)v_{t+1}^{\lambda}+\lambda V^{\lambda}_{t+1}&\text{if}~{}~{}t<H\\ v_{H}^{\lambda}&\text{if}~{}~{}t=H\end{cases},$ (2)
where $v_{t}^{\lambda}=\mathbb{E}[v_{\phi}(V_{t}^{\lambda}\mid s_{t})]$ denotes the expected value of the distribution predicted by the critic. It is trained using maximum likelihood by minimizing
$\mathcal{L}(v_{\phi})=\mathbb{E}_{p_{\theta},\pi_{\phi}}\left[-\sum_{t=1}^{H-1}\mathrm{ln}\ v_{\phi}(V_{t}^{\lambda}\mid s_{t})\right].$ (3)
The actor $\pi_{\phi}(a_{t}|s_{t})$ is then updated by maximizing the $\lambda$-return regularized by actor entropy $\mathbf{H}[a_{t}|s_{t}]$:
$\mathcal{L}(\pi_{\phi})$ $=\mathbb{E}_{p_{\theta},\pi_{\phi}}\left[\sum^{H-1}_{t=1}\left(-v_{t}^{\lambda}-\eta\cdot\mathbf{H}[a_{t}|s_{t}]\right)\right].$ (4)
In practice, we found that using the pre-trained world model alone fails to work well in many cases, especially for hard-exploration tasks. To identify the reasons, we need to rethink the role of each component in policy updates. Given the policy update rule in Eq. 4, the policy update is infected by the distribution of imagined trajectories $\tilde{\tau}(s,a)=p_{0}(s)\prod_{t=0}^{H-1}\pi_{\phi}(a_{t}|s_{t})p_{\theta}(s_{t+1}|s_{t},a_{t})$ and the reward model $r_{\theta}(s_{t},a_{t})$, i.e., the initial state distribution $p_{0}(s)$, the policy $\pi_{\phi}(a|s)$, the dynamics model $p_{\theta}(s_{t+1}|s_{t},a_{t})$, and the reward model $r_{\theta}(s_{t},a_{t})$. Intuitively, as shown in Fig. 3, the policy will learn a good policy if the imagined trajectories $\tilde{\tau}$ contain “promising” data for policy learning and the reward function $r_{\theta}(s_{t},a_{t})$ can correctly label those trajectories. The world model pre-training gives a better initialization of $p_{\theta}(s_{t+1}|s_{t},a_{t})$. However, when the online data distribution is narrow – as in hard exploration tasks – the initial state distribution $p_{0}(s)$ can also be narrow, which prevents the actor from reaching high-reward states during imagination. Also, the reward function $r_{\theta}(s,a)$ trained on the narrow online data distribution will incorrectly assign rewards for transitions generated by the world model pre-trained on a broader distribution. That is, the agent fails to improve the policy even if “promising” trajectories exist in $\tilde{\tau}$, since incorrect rewards are assigned. Furthermore, fine-tuning the world model $p_{\theta}(s_{t+1}|s_{t},a_{t})$ on the narrow distribution causes catastrophic forgetting. In the following, we improve the agents’ performance by reusing non-curated offline data for i) augmenting the initial state distribution $p_{0}(s)$, ii) collecting diverse online trajectories, and iii) improving the reward predictions. See Fig. 3 for an illustrative motivation.
Retrieval-based Experience Rehearsal
Figure 3: Motivation of using experience rehearsal and execution guidance. During fine-tuning, experience rehearsal augments the initial state $p_{0}(s)$ used for model rollout, which enables the model rollout to start from informative states. Execution guidance helps the agent to collect data that is close to the offline data distribution, which enriches the diversity of the collected data and improves the reward model’s accuracy on a broader state distribution.
Experience replay is a simple yet effective approach to prevent catastrophic forgetting (Khetarpal et al., 2022) as well as improving policy learning (Ball et al., 2023; Li et al., 2023). In the RL setting, Ball et al. (2023) demonstrates promising results by replaying offline data to improve sample efficiency but requires the offline data to be reward-labeled. ExPLORe (Li et al., 2023) reuses single-embodiment but reward-free offline data to encourage exploration by labeling offline data with UCB rewards. Unlike the replay scheme used in ExPLORe, where the offline datasets are relatively small and well-structured, directly replaying the non-curated offline data can be challenging. Our dataset can be 100$\sim$ 100$\times$ 100 × larger and consist of different tasks and embodiments. As such, directly labeling the offline dataset with UCB reward or even replaying the dataset can be infeasible.
To enable the usage of the non-curated offline data, we propose a simple retrieval-based approach to filter the trajectories that are close to the downstream tasks. Specifically, we retrieve a subset of trajectories from the non-curated offline data based on the neural feature distance between the online samples and the trajectories in the offline dataset. We calculate the distance as :
$\mathbf{D}=||\text{e}_{\theta}(o_{\text{on}})-\text{e}_{\theta}(o_{\text{off}})||_{2},$ (5)
where $\text{e}_{\theta}$ is a feature extractor and $o_{\text{on}}$ and $o_{\text{off}}$ are images from the initial observation of trajectories from the online buffer and the offline dataset, respectively. In practice, we found that neural representation matters for retrieval and the encoder learned via world model pre-training works better than the general visual model R3M (Nair et al., 2022).
To efficiently search over the whole dataset, we pre-build the key-value pairs between the trajectories ID and its neural features and use Faiss (Douze et al., 2024) for similarity search. Once the key-value database is built, the buffer retrieval procedure only takes seconds. The retrieved offline buffer, although reward-free, contains embodiment-specific data and is feasible to use. The retrieved offline data can be used to prevent catastrophic forgetting in the world model and to augment the initial state distribution $p_{0}(s)$ used for trajectory imagination.
Figure 4: WPT enables sample-efficient RL training. We include 6 tasks from DMControl and Meta-World with pixel inputs. The mean and the corresponding 95% confidence interval across 3 seeds are plotted. Full results on 72 tasks are included in App. D.
Execution Guidance via Prior Actors
The common practice in RL training is to initialize the replay buffer with a random policy and gradually collect new data by interacting with the environment using the training policy. When using a pre-trained world model, it is preferred to start the training from a distribution that is close to the offline data distribution. This has three benefits: First, offline data usually contains useful information for policy training, such as near-expert trajectories or a broader state-action distribution. Second, the distribution shift between the offline and online data may ruin the pre-trained weights, so bringing the initial distribution close to the offline data is a good strategy to prevent this. Thirdly, the reward function can get better predictions for states that are close to the offline data distribution. This improves policy learning as the initial states used to generate the imagined trajectories $\tilde{\tau}$ are augmented with offline samples.
To guide the online data collection such that it remains close to the offline data distribution, we simply train a prior actor $\pi_{\text{bc}}$ via behavior cloning on the retrieved buffer. During online data collection, we switch between the policy prior $\pi_{\text{bc}}$ and the RL agent $\pi_{\phi}$ according to a pre-defined schedule. Specifically, at the beginning of an episode, we decide whether to use $\pi_{\text{bc}}$ according to the scheduled probability. If the $\pi_{\text{bc}}$ is selected to use, we then randomly select the starting time step $t_{\text{bc}}$ and duration $H$ of using $\pi_{\text{bc}}$. $\pi_{\phi}$ is then used for the rest of the time steps. This procedure is similar to JSRL (Uchendu et al., 2023) but with a couple of differences: i) WPT leverages non-curated offline data while JSRL uses task-specific offline data and ii) WPT shows the superiority of a model-based approach for leveraging offline data while JSRL is a model-free approach. As a result, we show that WPT obtains better performance than JSRL on tested tasks. The full algorithm is presented in App. C.
4 Experiments
In this section, we investigate our WPT in two different RL settings. In the first setting, we focus on how the proposed WPT helps to improve the sample efficiency of RL training. In the second setting, we investigate whether WPT enables fast task adaptation, which is particularly important to empower the agent’s life-long learning ability. Lastly, we conduct an ablation study to investigate the role of proposed design choices.
4.1 Dataset
Our dataset consists of data from two different benchmarks: DeepMind Control Suite (DMControl) and Meta-World, visualized in Fig. 2 and Fig. 9 for the full list. For DMControl, we include 10k trajectories covering 5 embodiments collected by unsupervised RL agents (Pathak et al., 2017; Rajeswar et al., 2023), which are trained via curiosity without any task-related information. For Meta-World, we collect 50k trajectories spinning 50 tasks by executing the pre-trained RL agents from TDMPCv2 (Hansen et al., 2024) with injected Gaussian noise with $\sigma=[0.1,0.3,0.5,1.0,2.0]$ to mimic imperfect demonstrations. In total, the offline dataset consists of 60k trajectories with 10M state-action pairs from 6 different embodiments.
4.2 WPT Enables Sample-efficient Fine-tuning
Tasks   We evaluate our method on six selected pixel-based continuous control tasks from DMControl and Meta-World. The chosen tasks cover different challenges in RL. Except for learning from high-dimensional observations, we include a hard exploration version of Cheetah Run by setting zero rewards below a reward threshold and applying an action penalty. We include Walker Run and Quadruped Walk for complex dynamics. We further include challenging manipulation tasks from MetaWorld.
Baselines   In our experiments, we consider four different baselines for leveraging reward-free offline data. Note that as these baselines are not designed to handle multi-embodiment offline data, we preprocess the offline data to only contain task-related data for the baselines. Although this is an unfair comparison for WPT as baselines eliminate the difficulty raised with the large non-curated dataset, we still outperform these baselines by a large margin on tested tasks. The compared baselines are:
  • R3M pre-trains visual representation aiming to boost policy learning. We use a ResNet-18 backbone and train additional policy and value heads via RL. We include this baseline to compare with a representative approach for visual pre-training.
  • UDS suggests to label offline data with zero. UDS is an offline RL method, which cannot be directly used for comparison. We instead set zero reward for offline data and use RLPD (Ball et al., 2023) for policy training.
  • ExPLORe labels offline data with UCB reward. We follow the implementation from Li et al. (2023) but use ensembles for reward estimation, which has better performance.
  • JSRL-BC (Uchendu et al., 2023) collects online data with both the training policy and a prior policy trained on offline data via IL. We include JSRL-BC to compare with different ways of using offline data.
Figure 5: WPT enables fast task adaptation. We train an RL agent to control an Ant robot from DMControl to finish a series of tasks incrementally. WPT outperforms the widely used baseline by a decent margin by properly leveraging non-curated offline data.
Figure 6: Ablation study on the role of each component. “P” represents world model pretraining, “ER” means experience rehearsal, and “G” represents execution guidance. Together with the proposed retrieval-based experience rehearsal and execution guidance, world model pre-training boosts RL performance on a wide range of tasks.
Results   Figure 4 shows comparison results with baselines. Our method outperforms or matches all compared baselines by a large margin. Compared to R3M, WPT stresses the importance of world model pre-training as well as reusing offline data during fine-tuning compared to representation learning only. R3M fails to improve sample efficiency on most of the tested tasks. The same conclusion is obtained in Hansen et al. (2023). UDS and ExPLORe reuse offline data by labeling offline data with zero rewards and UCB rewards respectively. The labeled offline data is concatenated with online data for off-policy updates. We find UDS only shows slightly better performance on the Walker Run task compared to R3M and JSRL-BC, showing the ineffectiveness of directly labeling offline data with zero rewards. ExPLORe shows better performance on 2/3 locomotion tasks compared to other baselines and shows signs of meaningful progress on two challenging manipulation tasks. However, WPT still outperforms ExPLORe by a large margin. Compared to ExPLORe, WPT demonstrates the superiority of world model pre-training as well as execution guidance.
Compared to JSRL-BC, our method also demonstrates a clear performance boost. JSRL-BC’s performance heavily depends on the distribution of offline datasets. It can achieve good performance if a good prior actor can be extracted from the offline data. However, in our case, since we aim to lift the assumption of expert trajectories, the pre-trained BC agent usually fails to work well in the target task. As a result, JSRL-BC is only slightly better than the rest baselines on the Cheetah Run Hard, Assembly, and Shelf Place tasks while achieving similar results on the rest. In contrast, WPT works well on leveraging non-expert offline data. For example, the Quadruped Walk shows that WPT can nicely benefit from exploratory offline data collected by unsupervised RL, which enables it to control the Ant robot to walk forward from pixel inputs with only 100 trials.
In Fig. 1 and App. D, we further compare our method with two widely used learning-from-scratch baselines, Dreamerv3 and DrQv2, on 72 tasks. We show that by leveraging the non-curated offline data, WPT can clearly improve the sample efficiency on a wide range of tasks. Furthermore, our method achieves promising performance on several hard exploration tasks, while learning-from-scratch baselines fail.
4.3 WPT Enables Fast Task Adaptation
In many real-world applications, we want the robot to continuously adapt to new tasks. In this section, we investigate whether WPT benefits task adaptation. In this setting, the agent is required to incrementally solve a sequence of tasks. The setting is very similar to continual reinforcement learning (CRL) or life-long RL (Parisi et al., 2019; Khetarpal et al., 2022) but with a limited set of tasks for simplicity. Note that CRL has a broad scope; assumptions and experiment setups vary among methods, which makes it difficult to set up a fair comparison with other methods. Instead of proposing a SOTA CRL method, we aim to demonstrate that our method offers a simple yet general recipe to leverage previous data that also fits the CRL setting.
Setup & baselines   We set our continual multi-task adaption experiment based on the Ant robot from the DeepMind Control Suite. Specifically, the agent needs to sequentially solve stand, walk, run, jump, roll, and roll fast tasks. We train 300K environment steps for each task. To solve one task, the agent accesses all previous experiences as well as the model’s weights trained on previous tasks. To have a fair comparison, i.e., having comparable model parameters and eliminating the potential effects from pre-training on other tasks, we pre-train a small world model only on the Ant domain. We compare our method with PackNet (Mallya & Lazebnik, 2018), which is a common baseline for continual learning. PackNet iteratively prunes the actor’s parameters by keeping the weights with a larger magnitude while re-initializing the rest. By doing so, the actor maintains skills learned previously, which may help the learning of new tasks. For each new task, we fine-tune the actor model by iterative pruning while randomly initializing the critic model as rewards are not shared among tasks.
Results   Figure 5 shows policy performance on incremental task adaption. As we can see WPT outperforms baselines by a large margin. Compared to PackNet, WPT shows great potential for pre-training of world models with non-curated offline data together with experience rehearsal and actor prior. For the tested Ant domain, WPT enables the agent to adapt to new tasks within only 100 trials on all six tasks. With the limited sample budget, PackNet only achieves $\sim 20-60\%$ episodic returns of WPT. We would argue the diversity of the non-curated offline data contributes to WPT’s supreme performance on task adaptation.
4.4 Ablations
In this section, we investigate the role of each component of our method. We use the same set of tasks used in Fig. 4. As shown in Fig. 6, world model pre-training shows promising results when the offline dataset consists of diverse trajectories, such as data collected by exploratory agents, while fails to work well when the offline data distribution is relatively narrow as in the Meta-World tasks. This is due to i) world model pre-training alone failing to fully leverage the rich action prior in the offline data and ii) the distributional shift between offline and online data hurts the world model fine-tuning. The proposed retrieval-based experience rehearsal and execution guidance helps the agent benefit from action prior in the offline data and enriches the online sample diversity, which together enables WPT to achieve strong performance on a wide range of tasks.
5 Conclusion
We propose WPT, a simple yet efficient approach to leverage non-curated offline data. We show that the generalist world model pre-trained on non-curated data can boost RL training spanning multiple tasks and multiple embodiments. Together with retrieval-based experience rehearsal and execution guidance, WPT outperforms baselines on a wide range of tasks, including 22 locomotion tasks and 50 manipulation tasks. WPT unlocks ample sources of offline data and our results show that leveraging this non-curated data leads to strong performance. Nevertheless, our method can be improved in multiple ways, for example, by extending our methods to real-world applications as well as proposing novel world model architectures.
Impact Statement
This paper contributes to the field of reinforcement learning (RL), with potential applications including robotics and autonomous machines. While our methods hold promise for advancing technology, they could also be applied in ways that raise ethical concerns, such as in autonomous machines exploring the world and making decisions on their own. However, the specific societal impacts of our work are broad and varied, and we believe a detailed discussion of potential negative uses is beyond the scope of this paper. We encourage a broader dialogue on the ethical use of RL technology and its regulation to prevent misuse.
Appendices
Appendix A Implementation Details
A.1 Behavior Cloning
The Behavior Cloning methods used in both the execution guidance of WPT and JSRL-BC are the same. We use a four-layer convolutional neural network (LeCun et al., 1995) with kernel depth [32, 64, 128, 256] following a three-layer MLPs with LayerNorm (Ba et al., 2016) after all linear layers.
We list the used encoder and actor architectures for reference.
1class Encoder(nn.Module):
2 def __init__(self, obs_shape):
3 super().__init__()
4 assert obs_shape == (9, 64, 64), f’obs_shape is {(obs_shape)}, but expect (9, 64, 64)’ # inputs shape
5
6 self.repr_dim = (32 * 8) * 2 * 2
7 _input_channel = 9
8
9 self.convnet = nn.Sequential(
10 nn.Conv2d(_input_channel, 32, 4, stride=2), # [B, 32, 31, 31]
11 nn.ELU(),
12 nn.Conv2d(32, 32*2, 4, stride=2), #[B, 64, 14, 14]
13 nn.ELU(),
14 nn.Conv2d(32*2, 32*4, 4, stride=2), #[B, 128, 6, 6]
15 nn.ELU(),
16 nn.Conv2d(32*4, 32*8, 4, stride=2), #[B, 256, 2, 2]
17 nn.ELU())
18 self.apply(utils.weight_init)
19
20 def forward(self, obs):
21 B, C, H, W = obs.shape
22
23 obs = obs / 255.0 - 0.5
24 h = self.convnet(obs)
25 # reshape to [B, -1]
26 h = h.view(B, -1)
27 return h
1class Actor(nn.Module):
2 def __init__(self, repr_dim, action_shape, feature_dim=50, hidden_dim=1024):
3 super().__init__()
4 self.trunk = nn.Sequential(nn.Linear(repr_dim, feature_dim),
5 nn.LayerNorm(feature_dim), nn.Tanh())
6
7 self.policy = nn.Sequential(nn.Linear(feature_dim, hidden_dim),
8 nn.LayerNorm(hidden_dim), nn.ELU(),
9 nn.Linear(hidden_dim, hidden_dim),
10 nn.LayerNorm(hidden_dim), nn.ELU(),
11 nn.Linear(hidden_dim, action_shape[0]))
12
13 self.apply(utils.weight_init)
14
15 def forward(self, obs, std):
16 h = self.trunk(obs)
17 reutrn self.policy(h)
A.2 JSRL+BC
Jump-start RL (Uchendu et al., 2023) is proposed as an offline-to-online RL method. It includes two policies, a prior policy $\pi_{\theta_{1}}(a|s)$ and a behavior policy $\pi_{\theta_{2}}(a|s)$, where the prior policy is trained via offline RL methods and the behavior policy is updated during the online learning stage. However, offline RL requires the offline dataset to include rewards for the target task. To extract behavior policy from the offline dataset, we use the BC agent described above as the prior policy. During online training, in each episode, we randomly sample the rollout horizon $h$ of the prior policy from a pre-defined array np.arange(0, 101, 10). We then execute the prior policy for $h$ steps and switch to the behavior policy until the end of an episode.
A.3 ExPLORe
For the ExPLORe baseline, we follow the original training code   [^1] Source code of ExPLORe https://github.com/facebookresearch/ExPLORe. We sweep over several design choices: i) kernel size of the linear layer used in the RND and reward models: [256 (default), 512]; ii) initial temperature value: [0.1 (default), 1.0]; iii) whether to use LayerNorm Layer (no by default); iv) learning rate: [1e-4, 3e-4 (default)]. However, we fail to obtain satisfactory performance. There are several potential reasons: i) the parameters used in the ExPLORe paper are tuned specifically to their setting, where manipulation tasks and near-expert trajectories are used; ii) the coefficient term of the RND value needs to be tuned carefully for different tasks and the reward should also be properly normalized.
To achieve reasonable performance and eliminate the performance gap caused by implementation-level details, we make the following modifications: i) we replace the RND module with ensembles to calculate uncertainty; ii) the reward function share the latent space with actor and critic.
Appendix B More Ablation Study
We include per-task results used in Sec. 4.4 in Fig. 7.
Figure 7: Ablation study of the role of each component.
We further compare results using uncertainty-based rewards as suggested in (Li et al., 2023) compared with our expectation guidance. Instead of using UCB for reward labels, we use Optimistic Thompson Sampling (OTS) (Hu et al., 2023b), which is less sensitive w.r.t the selection of hyperparameters. As shown in Fig. 8, our method outperforms the variant using OTS on Assembly and Stick Pull by a large margin, showing the effectiveness of using execution guidance.
Figure 8: Ablation study of using uncertainty-based reward labeling.
Appendix C Algorithm
The full algorithm is described in Alg. 1.
Algorithm 1 Generalist World Model Pre-Training for Efficient RL
0:  Non-curated offline data $\mathcal{D}_{\text{off}}$, Online data $\mathcal{D}_{\text{on}}\leftarrow\emptyset$, Retrieval data $\mathcal{D}_{\text{retrieval}}\leftarrow\emptyset$              World model $f_{\theta},q_{\theta},p_{\theta},d_{\theta}$              Policy $\pi_{\phi_{\text{RL}}}$, $\pi_{\phi_{\text{BC}}}$, Value function $v_{\phi}$ and Reward $r_{\xi}$.
   // Task-Agnostic World Model Pre-Training
  for num. pre-train steps do
     Randomly sample mini-batch $\mathcal{B_{\text{off}}}:\{o_{t},a_{t},o_{t+1}\}_{t=0}^{T}$ from $\mathcal{D}_{\text{off}}$.
     Update world model $f_{\theta},q_{\theta},p_{\theta},d_{\theta}$ by minimizing Eq. 1 on sampled batch $\mathcal{B}$.
  end for
   // Task-Specific Training
   // Experience Retrieval
  Collect one initial observation $o^{0}_{\text{on}}$ from the environment.
  Compute the visual similarity between $o_{\text{on}}$ and initial observations of trajectories $o_{\text{off}}$ in $\mathcal{D}_{\text{off}}$ using Eq. 5.
  Select R trajectories according to Eq. 5 and fill $\mathcal{D}_{\text{retrieval}}$.
   // Behavior Cloning Policy Training
  for num. bc updates do
     Randomly sample mini-batch $\mathcal{B_{\text{retrieval}}}:\{o_{i},a_{i}\}_{i=0}^{N}$ from $\mathcal{D}_{\text{retrieval}}$.
     Update $\pi_{\phi_{\text{BC}}}$ by minimizing $-\frac{1}{N}\sum_{i=0}^{N}\log\pi_{\phi_{\text{BC}}}(a_{t}|o_{t})$.
  end for
   // Task-Specific RL Fine-Tuning
  for num. episodes do
      // Collect Data
     Decide whether to use $\pi_{\phi_{\text{BC}}}$ according to the predefined schedule.
     if Select $\pi_{\phi_{\text{BC}}}$ then
        Randomly select the starting time step $k$ and the rollout horizon $H$.
     end if
     $t\leftarrow 0$
     while  $t\leq$ episode length do
        $a_{t}=\pi_{\phi_{BC}}(a_{t}|o_{t})$ if Use $\pi_{\phi_{\text{BC}}}$ and $k\leq t\leq H$ else $a_{t}=\pi_{\phi_{RL}}(a_{t}|o_{t})$.
        Interact the environment with $a_{t}$. Store $\{o_{t},a_{t},r_{t},o_{t+1}\}$ to $\mathcal{D}_{\text{on}}$.
        $t\leftarrow t+1$
     end while
      // Update Models
     for num. grad steps do
        Randomly sample mini-batch $\mathcal{B}_{\text{on}}:\{o_{t},a_{t},r_{t},o_{t+1}\}_{t=0}^{T}$ from $\mathcal{D}_{on}$ and $\mathcal{B}_{\text{retrieval}}:\{o_{t},a_{t},r_{t},o_{t+1}\}_{t=0}^{T}$ from $\mathcal{D}_{\text{retrieval}}$.
        Update world model $f_{\theta},q_{\theta},p_{\theta},d_{\theta}$ by minimizing Eq. 1 on sampled batch $\{\mathcal{B_{\text{on}},\mathcal{B}_{\text{retrieval}}}\}$.
[…]
         // Update policy and value function
        Generate imaginary trajectories $\tilde{\tau}=\{s_{t},a_{t},s_{t+1}\}_{t=0}^{T}$ by rolling out $h_{\theta},p_{\theta}$ with policy $\pi_{\phi_{\text{RL}}}$.
        Update policy $\pi_{\phi_{\text{RL}}}$ and value function $v_{\phi}$ with Eq. 4.
     end for
  end for
Appendix D Full Results
In Table 2 and Table 3, we list the success rate of 50 Meta-World benchmark tasks with pixel inputs. In Table 4, we list the episodic return of DMControl of 22 tasks. We compare WPT at 150k samples with two widely used baselines Dreamer v3 and DrQ v2 at both 150k samples and 1M samples. We mark the best result with a bold front at 150k samples and use underlining to mark the highest score overall.
D.1 Meta-World Benchmark
Table 2: Success rate of Meta-World benchmark with pixel inputs.
Tasks Dreamer v3 @ 1M DrQ v2 @ 1M Dreamer v3 @ 150k DrQ v2 @ 150k WPT (ours) @ 150k
Assembly 0.0 0.0 0.0 0.0 0.2
Basketball 0.0 0.97 0.0 0.0 0.4
Bin Picking 0.0 0.93 0.0 0.33 0.8
Box Close 0.13 0.9 0.0 0.0 0.9
Button Press 1.0 0.7 0.47 0.13 0.9
Button Press Topdown 1.0 1.0 0.33 0.17 1.0
Button Press Topdown Wall 1.0 1.0 0.73 0.63 1.0
Button Press Wall 1.0 1.0 0.93 0.77 1.0
Coffee Button 1.0 1.0 1.0 1.0 1.0
Coffee Pull 0.6 0.8 0.0 0.6 0.6
Coffee Push 0.67 0.77 0.13 0.2 0.7
Dial Turn 0.67 0.43 0.13 0.17 0.67
Disassemble 0.0 0.0 0.0 0.0 0.0
Door Close - - - - 1.0
Door Lock 1.0 0.93 0.6 0.97 0.9
Door Open 1.0 0.97 0.0 0.0 0.8
Door Unlock 1.0 1.0 1.0 0.63 0.8
Drawer Close 0.93 1.0 0.93 1.0 0.9
Drawer Open 0.67 0.33 0.13 0.33 1.0
Faucet Open 1.0 1.0 0.47 0.33 1.0
Faucet Close 0.87 1.0 1.0 1.0 0.8
Hammer 1.0 1.0 0.07 0.4 1.0
Hand Insert 0.07 0.57 0.0 0.1 0.4
Handle Press Side 1.0 1.0 1.0 1.0 1.0
Handle Press 1.0 1.0 0.93 0.97 1.0
Handle Pull Side 0.67 1.0 0.67 0.6 1.0
Handle Pull 0.67 0.6 0.33 0.6 1.0
Lever Pull 0.73 0.83 0.0 0.33 0.8
More results in the next page, see Table 3
Table 3: Success rate of Meta-World benchmark with pixel inputs (Cont.).
Tasks Dreamer v3 @ 1M DrQ v2 @ 1M Dreamer v3 @ 150k DrQ v2 @ 150k WPT (ours) @ 150k
Peg Insert Side 1.0 1.0 0.0 0.27 1.0
Peg Unplug Side 0.93 0.9 0.53 0.5 0.8
Pick Out of Hole 0.0 0.27 0.0 0.0 0.3
Pick Place Wall 0.2 0.17 0.0 0.0 0.5
Pick Place 0.67 0.67 0.0 0.0 0.2
Plate Slide Back Side 1.0 1.0 0.93 1.0 1.0
Plate Slide Back 1.0 1.0 0.8 0.97 1.0
Plate Slide Side 1.0 0.9 0.73 0.5 0.5
Plate Slide 1.0 1.0 0.93 1.0 1.0
Push Back 0.33 0.33 0.0 0.0 0.2
Push Wall 0.33 0.57 0.0 0.0 0.9
Push 0.26 0.93 0.0 0.13 0.7
Reach 0.87 0.73 0.67 0.43 0.3
Reach Wall 1.0 0.87 0.53 0.7 0.9
Shelf Place 0.4 0.43 0.0 0.0 0.87
Soccer 0.6 0.3 0.13 0.13 0.67
Stick Push 0.0 0.07 0.0 0.0 0.4
Stick Pull 0.0 0.33 0.0 0.0 0.67
Sweep Into 0.87 1.0 0.0 0.87 0.9
Sweep 0.0 0.73 0.0 0.3 0.6
Window Close 1.0 1.0 0.93 1.0 0.8
Window Open 1.0 0.97 0.6 1.0 0.9
Mean 0.900 0.753 0.360 0.442 0.750
Medium 0.870 0.900 0.130 0.330 0.800
D.2 DMControl Benchmark
Table 4: Episodic return of DMControl benchmark with pixel inputs.
Tasks Dreamer v3 @ 500k DrQ v2 @ 500k Dreamer v3 @ 150k DrQ v2 @ 150k WPT (ours) @ 150k
CartPole Balance 994.3 992.3 955.8 983.3 995.0
Acrobot Swingup 222.1 30.3 85.2 20.8 50.3
Acrobot Swingup Sparse 2.5 1.17 1.7 1.5 26.9
Acrobot Swingup Hard -0.2 0.3 2.0 0.4 10.7
Walker Stand 965.7 947.6 946.2 742.9 969.4
Walker Walk 949.2 797.8 808.9 280.1 959.1
Walker Run 616.6 299.3 224.4 143.0 728.0
Walker Backflip 293.6 96.7 128.2 91.7 306.0
Walker Walk Backward 942.9 744.3 625.9 470.9 863.6
Walker Walk Hard -2.1 -9.5 -4.7 -17.1 878.3
Walker Run Backward 363.8 246.0 229.4 167.4 349.3
Cheetah Run 843.7 338.1 621.4 251.2 526.1
Cheetah Run Front 473.8 202.4 143.1 108.4 360.5
Cheetah Run Back 657.4 294.4 407.6 171.2 446.0
Cheetah Run Backwards 693.8 384.3 626.6 335.6 542.2
Cheetah Jump 597.0 535.6 200.8 251.8 634.1
Quadruped Walk 369.3 258.1 145.2 76.5 933.6
Quadruped Stand 746.0 442.2 227.2 318.9 936.4
Quadruped Run 328.1 296.5 183.0 102.8 802.5
Quadruped Jump 689.6 478.3 168.3 190.5 813.5
Quadruped Roll 663.9 446.0 207.9 126.2 970.8
Quadruped Roll Fast 508.8 366.9 124.8 164.7 782.0
Mean 541.81 372.23 320.86 226.49 631.10
Medium 606.8 318.70 204.35 166.05 755.0
Appendix E Hyperparameters
In this section, we list important hyperparameters used in WPT.
Table 5: Hyperparameters used in WPT.
Hyperparameter Value
Pre-training
Stacked images 1
Pretrain steps 200,000
Batch size 16
Sequence length 64
Replay buffer capacity Unlimited
Replay sampling strategy Uniform
RSSM
     Hidden dimension 12288
     Deterministic dimension 1536
     Stochastic dimension 32 * 96
     Block number 8
     Layer Norm True
CNN channels [96, 192, 384, 768]
Activation function SiLU
Optimizer
     Optimizer Adam
     Learning rate 1e-4
     Weight decay 1e-6
     Eps 1e-5
     Gradient clip 100
Fine-tuning
Warm-up frames 15000
Execution Guidance Schedule linear(1,0,50000) for DMControl
linear(1,0,1,150000) for Meta-Wolrd
Action repeat 2
Offline data mix ratio 0.25
Discount 0.99
Discount lambda 0.95
MLPs [512, 512, 512]
MLPs activation SiLU
Actor critic learning rate 8e-5
Actor entropy coef 1e-4
Target critic update fraction 0.02
Imagine horizon 16
Appendix F Task Visualization
Figure 9: Visualization of tasks from DMControl and Meta-World used in our paper.